• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name A. Puckett

    Log entry time 08:02:19 on April24,2007

    Entry number 199831

    keyword=Owl shift summary

    At the beginning of the shift, the HARP scans were completed, but a raster FSD card misbehaving kept us down for ~1 hour. MCC was able to fix this after summoning the PSS-on-call, expected they would need to access Hall A to fix, but ultimately did not. It is possible that the raster size being set to "0.2x0.2" was causing the hiccup. By about 2:11, the raster had been fixed, and by 3:00, we had the raster size and beam position to spec, and then we moved to 4He, and we have been running on 4He ever since, at 90 muA. Beam has been fairly trippy.